Consider the following hypotheses:

H0: μ = 410
HA: μ ≠ 410

The population is normally distributed with a population standard deviation of 46. (You may find it useful to reference the appropriate table: z table or t table)

a-1. Calculate the value of the test statistic with x−x− = 421 and n = 85. (Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)



a-2. What is the conclusion at the 10% significance level?

  • Do not reject H0 since the p-value is greater than the significance level.

  • Do not reject H0 since the p-value is less than the significance level.

  • Reject H0 since the p-value is greater than the significance level.

  • Reject H0 since the p-value is less than the significance level.



a-3. Interpret the results at αα = 0.10.

  • We cannot conclude that the population mean differs from 410.

  • We conclude that the population mean differs from 410.

  • We cannot conclude that the sample mean differs from 410.

  • We conclude that the sample mean differs from 410.



b-1. Calculate the value of the test statistic with x−x− = 397 and n = 85. (Negative value should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)



b-2. What is the conclusion at the 5% significance level?

  • Reject H0 since the p-value is greater than the significance level.

  • Reject H0 since the p-value is less than the significance level.

  • Do not reject H0 since the p-value is greater than the significance level.

  • Do not reject H0 since the p-value is less than the significance level.



b-3. Interpret the results at αα = 0.05.

  • We conclude that the population mean differs from 410.

  • We cannot conclude that the population mean differs from 410.

  • We conclude that the sample mean differs from 410.

  • We cannot conclude that the sample mean differs from 410.
Answer #1

Test statistic,
z = (xbar - mu)/(sigma/sqrt(n))
z = (421 - 410)/(46/sqrt(85))
z = 2.2

P-value Approach
P-value = 0.0278
As P-value < 0.1, reject the null hypothesis.

Reject H0 since the p-value is less than the significance level.

We conclude that the sample mean differs from 410.

#2.
Test statistic,
z = (xbar - mu)/(sigma/sqrt(n))
z = (397 - 410)/(46/sqrt(85))
z = -2.61

P-value Approach
P-value = 0.0091

Reject H0 since the p-value is less than the significance level.

We conclude that the population mean differs from 410.